Peterson keeps both cows and chickens in his boma. if there are 13 animals and altogether they have 34 legs, how many cows and how many chickens does he have..?
Answer by ikleyn(52776)

.

                From Wikipedia, the free encyclopedia (https://en.wikipedia.org/wiki/Boma_(enclosure) )

    A boma is a livestock enclosure, community enclosure, stockade, corral, small fort or a district government 
    office and community used in many parts of the African Great Lakes region, as well as Central and Southern Africa. 
    It is particularly associated with community decision making. It is incorporated into many African languages, 
    as well as colonial varieties of English, French and German.


                         Solution


Let x be the number of cows.

Then the number of chicken is (13-x).


There is no need to remind you that the cows have 4 legs each, while the chicks have only 2 legs.


So, you write this equation for legs


    4x + 2(13-x) = 34


Simplify and find x


    4x + 26 - 2x = 34

    4x - 2x = 34 - 26

       2x   = 8

        x   = 8/2 = 4.


ANSWER.  4 cows and (13-4) = 9 chicken.


CHECK.   4*4 + 9*2 = 16 + 18 = 34,  total legs.    ! Correct !

Solved.

---------------

To see similar  (and different)  solved problems,  look into the lessons
    - Problem on two-wheel and three-wheel bicycles
    - Problem on animals at a farm
    - Problem on pills in containers
    - What type of problems are these?
in this site.

You will find there different methods of solution - from one equation to two equations, and even without using any equations.


///////////////


Do not forget to post your  "THANKS"  to me for my teaching.